# Detecting the hidden population of low-mass haloes in strong lenses

### v1

@@ Abstract
A generic prediction of particle dark matter theories is that a large population of dark matter substructures should reside inside the host haloes of galaxies. In gravitational imaging, strong gravitational lens observations are used to detect individual objects from this population, if they are large enough to perturb the strongly lensed images. We show here that low-mass haloes, below the individually detectable mass limit, have a detectable effect on the lensed images when in large numbers, which is the case in cold dark matter (CDM). We find that, in CDM, this population causes an excess of 40 per cent in the number of detected subhaloes for HST-like strong lens observations. We propose a pseudo-mass function to describe this population, and fit for its parameters from the detection data. We find that it mostly consists of objects two orders of magnitude in mass below the detection limit of individual objects. We show that including this modification, so that the effect of the population is correctly predicted, can dramatically improve the available constraints on dark matter from strong lens observations. We repeat our experiments using models that contain varying amounts of angular structure in the lens galaxy. We find that these multipole perturbations are strongly degenerate with the population signal, to an even greater extent than with individual subhaloes. This further highlights the need for better understanding of the angular mass structure of lens galaxies, so that the maximum information can be extracted from strong lens observations for dark matter inference.
@@
pubyear: 2025pagerange: Detecting the hidden population of low-mass haloes in strong lensesReferences
@@ 1 Introduction
In the gravitational imaging method, the lens galaxy is first modelled with some mass distribution which is smooth on the scale of the lens galaxy. Then, free-form, pixellated corrections to the lensing potential are found which improve the fit, with some regularisation condition to impose smoothness (Vegetti & Koopmans, 2009). The density corrections corresponding to the corrected potential are inpsected. If a dark perturbing object exists in the field of view, it appears as a concentrated overdensity in the pixellated grid (Vegetti et al., 2012; Cao et al., 2025). Its properties can then be measured by fitting parametric models, comparing with the free-form corrected density map. The significant advantage of gravitational imaging is its agnosticism towards the specific form of the perturbation. No assumption needs to be made on the concentration of the perturbing object, its mass profile, whether or not it is at the redshift of the lens, or the ultimate number of perturbing objects in the field of view. If the data favour such concentrated overdensities, they will appear in the corrected convergence map without any necessary assumptions from the investigator on their properties.
@@
To overcome this issue, especially in the context of large numbers of new strong lenses becoming available from e.g. Euclid, forward-modelling methods have been proposed (e.g. Brehmer et al., 2019; Gilman et al., 2020; Anau Montel et al., 2023; Wagner-Carena et al., 2024; Filipp et al., 2024; Coogan et al., 2024). These methods treat the specifics of individual detections essentially as nuissance parameters, and instead seek to infer the dark matter properties directly. Typically, strong lens data are compared with large amounts of simulated data, simulated according to different dark matter models with the entire predicted population of low-mass haloes injected into the image. The likelihood of each dark matter model is then directly computed via comparison with the data. These methods fall in the category often described as approximate Bayesian computation or simulation based inference.
An obvious advantage is that the dark matter model is directly inferred, without the need to explicitly and expensively calculate the sensitivity. But this inference can be relied upon only as far as the assumptions made in producing the simulated data. The simulated data must account for all possible systematics at creation time, and there is no capacity to learn from the data after the fact, as with the potential corrections of gravitational imaging. This is also true for methods which forward model individual perturbers without gravitational imaging. In some cases these have been in good agreement with gravitational imaging (Nightingale et al., 2024). Other cases have produced false positive results, demonstrating the inherent risk of such a method (Hezaveh et al., 2016; Stacey et al., 2025).
However, forward-modelling methods have another significant advantage in that they can account for the effect of the entire population of perturbers, not just those large enough to be individually detectable with gravitational imaging. After all, in CDM, we expect hundreds or thousands of subhaloes larger than $10^{6}M_{\odot}$ to populate the haloes of lens galaxies. In gravitational imaging, the effect of this population of objects on the lensed images, if it exists, is not accounted for. A correctly implemented forward-modelling method could detect or rule out such an effect, resulting in a more powerful inference on the dark matter model than considering individually detected objects only. Many of the works cited above include this effect implicitly, but do not show systematically that it exists, or measure its properties.
@@ 2 Method
We use the machine learning-based sensitivity mapping and substructure detection method introduced in O’Riordan et al. (2023, hereafter O’R23) and extended in O’Riordan & Vegetti (2024, hereafter O’R24). Our results are based on applying these models to different realisations of the same $100$ lens systems used in O’R24, which were selected to resemble SLACS and BELLS Einstein ring systems previously studied for substructure.
@@ 2.2 Model training
To train the machine learning models we simulate strong lens observations at HST-like resolution and depth. The lens galaxies are power-law ellipsoids and the source surface brightness distributions are taken from a de-noised sample of Hubble Deep Field galaxy observations. The light profile of the main lens galaxy is subtracted, up to its remaining Poisson noise. Different sources are used in training and testing to avoid overfitting.
@@
As in O’R24, we account for non-ellipticity in the lens mass distribution using multipole components. These are Fourier-like perturbations to the main lens density profile. To examine their effect on substructure detection, three versions of the model are trained. $\mathcal{M}_{1}$ includes no multipole perturbations in the training data. $\mathcal{M}_{2}$ includes them only up to amplitudes of $1$ per cent, and $\mathcal{M}_{3}$ includes them up to 3 per cent. $\mathcal{M}_{1}$ essentially represents most strong lensing models used until recently, a power-law ellipsoid plus external shear. $\mathcal{M}_{2}$ represents the best extension to $\mathcal{M}_{1}$, with non-elliptical components informed by the isophotes of elliptical galaxies, a model which is being used more commonly at present. $\mathcal{M}_{3}$ represents an extreme extension of $\mathcal{M}_{2}$, so that effects due to multipoles can be more easily borne out.
@@ 2.3 Sensitivity mapping
To create a sensitivity map, the models are shown realisations of the same strong lens system with a test mass substructure that moves across all pixels and over a range of masses. Each realised image is passed through the model and the probability of the positive class, i.e., that some amount of substructure is present, is recorded. A probability threshold is chosen, in this case $5\sigma$. The sensitivity map itself then follows in a straightforward way, as simply the lowest mass of subhalo in each pixel where the detection probability reached the threshold. We name this mass $m_{\mathrm{map}}$. We compute a sensitivity map for each lens in this way using a test mass range of $10^{7}<M_{\mathrm{max}}<10^{11}$, with 21 mass steps spaced log-uniformly.
@@
Figure 3: Method to test the detectability of populations of individually undetectable substructures. A subhalo mass function, effectively the dark matter model, is chosen first, parametrised by $M_{\mathrm{hm}}$ and $f_{\mathrm{sub}}$. From this populations of subhaloes are drawn to produce images. The images are run through a subhalo detector CNN, which is the same used to create the sensitivity map for each lens. The sensitivity map and the subhalo mass function give $\mu_{\mathrm{sub}}$, the mean number of detectable objects, and in turn $p_{\mathrm{det}}$, the probability that the CNN makes a detection. The actual number of detections made by the CNN is $f_{\mathrm{det}}$. Comparing these quantitiies is the basis for the results in this paper.
@@ 2.4 Detecting a population of substructures
The number of detectable objects is Poisson distributed with mean $\mu_{\mathrm{sub}}$, so the probability that a given lens will have any detection, accouting for multiple individually detectable subhaloes, is
@@ 3.2 Excess as a function of subhalo mass
In region $(2)$, at $m_{\mathrm{cutoff}}\sim 10^{9.5}M_{\odot}$, the two curves begin to diverge. Figure˜2 shows that most pixels have no sensitivity below this mass. Accordingly, the predicted frequency of detections from the sensitivity map and mass fucntion becomes flat as all newly added subhaloes are too small to be individually detected. However, the actual frequency of detections by the CNN increases until $m_{\mathrm{cutoff}}\sim 10^{7}M_{\odot}$. In regime $(3)$, below $m_{\mathrm{cutoff}}\sim 10^{7}M_{\odot}$, the excess of detections stops increasing. This shows that subhaloes in the intermediate mass range, i.e., in region $(2)$, have a detectable effect on the lensed images. Subhaloes at the lowest mass, i.e., for $m_{\mathrm{cutoff}}\lesssim 10^{7}M_{\odot}$ and below, do not have a detectable effect.
Figure 5: Properties of the subhaloes in the detectable population for different values of $M_{\mathrm{hm}}$. Upper frame: the modified mass function in Eq.˜7 using the best fit values for our HST strong lenses. Lower frame: the ratio of the individual masses of the detectable population objects to the pixel sensitiviies in our data, randomly drawn.
@@ 3.3 Accounting for the detectable population
The number $\mu_{\mathrm{pop}}$ can be thought of as the number of extra detectable objects if all detectable ‘populations’ were converted to equivalent individual objects. We calculate $\mu_{\mathrm{pop}}$ using a modified subhalo mass function which artifically increases the number of objects of mass $m$ by a factor $g(m)$, that is,
@@
This introduces three new parameters. The normalisation $n_{\mathrm{pop}}$ represents the strength of the effect in causing detections, and may be roughly interpreted as the ratio of the number of subhaloes in a detectable population to an equivalant individually detectable object. The mean and variance of $g(m)$ are $m_{\mathrm{pop}}$ and $\sigma_{\mathrm{pop}}^{2}$, but these are only weak indicators of the mass range of the detectable population. This is because of the shape of the mass function, with which $g(m)$ will be multiplied, where every decade lower in mass brings a decade more in number (see again Fig.˜1). For example, if we find $m_{\mathrm{pop}}=10^{X}M_{\odot}$ with $\sigma_{\mathrm{pop}}=1$ (one order of magnitude in mass), much more than half of the population members will have mass $M_{\mathrm{max}}<10^{X}$.
@@
In the lower frame of Fig.˜5 we illustrate the size of the population members compared to typical detectable individual subhaloes. We draw subhaloes from the pseudo-mass function used in Eq.˜7 only and compare their masses with the sensitivity in a randomly drawn pixel from the $100$ mock lenses. For all values of $M_{\mathrm{hm}}$, the distribution peaks at $\sim 2$ orders of magnitude below the individually detectable mass. The absolute masses of this population (upper frame) depend on the data, so the peak in the detectable population mass function between $10^{8}M_{\odot}$ and $10^{9}M_{\odot}$ is specific to the BELLS and SLACS-like HST observations used here. However, it is very likely that the relative masses of the detecable population (lower frame) do not change with angular resolution. This is because sensitivity scales straightforwardly with angular resolution (Despali et al., 2022). Extrapolating the results here to higher angular resolutions would suggest that strong lens observations with e.g. VLBI, where objects as small as $10^{6}M_{\odot}$ are detectable, could be sensitive to populations of subhaloes whose constituent masses are $10^{5}M_{\odot}$ or lower. Replicating the results in this paper at that angular resolution is currently computationally intractable.
@@ 3.4 Improved constraints on dark matter
Here we consider the strength of constraint on the dark matter model, parametrised only by $M_{\mathrm{hm}}$, when the effect of the population either is or is not accounted for. We consider two simple instructive examples. In the first case, zero detections are made in our set of 100 lenses, which we label $\mathcal{D}_{0}$, and in the second case, 10 detections in 100 lenses, which we label $\mathcal{D}_{10}$. The case of 10 detections in 100 lenses represents an order of magniude estimate of a realistic scenario given current substructure search efforts and the quality of the available data.
@@ 3.5 Angular structure in the lens galaxy
Using the same models from O’R24 we repeat the experiment shown in Fig.˜4. This time, we use the two models described in Section˜2.2. These models were trained with mutlipole perturbations in the lensing galaxy. This makes their substructure detections, and sensitivity maps, much more conservative.
@@ 4 Discussion
Here we discuss some limitations and implications of our work. We do not discuss aspects related to machine learning specific issues, e.g. quality of the training data, model architecture, etc. These issues were discussed for the same models used here in O’R23 and O’R24. In O’R24 a detailed comparison was made between the sensitivity maps produced by our machine learning models, and classical sensitivity maps produced by Despali et al. (2022). We found the overall structure of the sensitivity maps to be the same. That is, for a given lens and source configuration, the topology of the sensitivity maps are similar. However, we find in general that the machine learning method produces more conservative sensitivity maps. This is for two reasons. First, its much larger ‘prior’, i.e., its training data, includes all possible source and lens configurations, rather than the combinations which closely resemble the lens being modelled, as in the classical method. Second, and more importantly, is the extreme difficulty of the problem as presented to it, where almost all of the subhaloes presented in the trainining data and labelled as positive cases, are not detectable because they are too small or too far from the lensed images, or not concentrated enough. The result is an extremely cautious model where a $10\sigma$ detection in the classical method may be more like a $5\sigma$ detection in our case. In the context of this work, this conservative approach is actually an advantage, as it means the machine learning models never make spurious false positive detections unless a genuinely detectable subhalo signal is present in the data.
4.1 Effect of lens galaxy angular structure

### v2

@@ Abstract
A generic prediction of particle dark matter theories is that a large population of dark matter substructures should reside inside the host haloes of galaxies. In gravitational imaging, strong gravitational lens observations are used to detect individual objects from this population, if they are large enough to perturb the strongly lensed images. We show here that low-mass haloes, below the individually detectable mass limit, have a detectable effect on the lensed images when in large numbers, which is the case in cold dark matter (CDM). We find that, in CDM, this population causes an excess of 40 per cent in the number of detected subhaloes for HST-like strong lens observations. We propose a pseudo-mass function to describe this population, and fit for its parameters from the detection data. We find that it mostly consists of objects two orders of magnitude in mass below the detection limit of individual objects. We show that including this modification, so that the effect of the population is correctly predicted, can improve the available constraints on dark matter from strong lens observations. We repeat our experiments using models that contain varying amounts of angular structure in the lens galaxy. We find that these multipole perturbations are degenerate with the population signal. This further highlights the need for better understanding of the angular mass structure of lens galaxies, so that the maximum information can be extracted from strong lens observations for dark matter inference.
@@
pubyear: 2026pagerange: Detecting the hidden population of low-mass haloes in strong lensesReferences
@@ 1 Introduction
In the gravitational imaging method, the lens galaxy is first modelled with some mass distribution which is smooth on the scale of the lens galaxy. Then, free-form, pixellated corrections to the lensing potential are found which improve the fit, with some regularisation condition to impose smoothness (Vegetti & Koopmans, 2009). The density corrections corresponding to the corrected potential are inspected. If a dark perturbing object exists in the field of view, it appears as a concentrated overdensity in the pixellated grid (Vegetti et al., 2012; Cao et al., 2025). Its properties can then be measured by fitting parametric models, comparing with the free-form corrected density map. The significant advantage of gravitational imaging is its agnosticism towards the specific form of the perturbation. No assumption needs to be made on the concentration of the perturbing object, its mass profile, or the ultimate number of perturbing objects in the field of view. If the data favour such concentrated overdensities, they will appear in the corrected convergence map without any necessary assumptions from the investigator on their properties.
@@
To overcome this issue, especially in the context of large numbers of new strong lenses becoming available from e.g. Euclid, forward-modelling methods have been proposed (e.g. Brehmer et al., 2019; Gilman et al., 2020; Anau Montel et al., 2023; Wagner-Carena et al., 2024; Filipp et al., 2024; Coogan et al., 2024). These methods treat the specifics of individual detections essentially as nuisance parameters, and instead seek to infer the dark matter properties directly. Typically, strong lens data are compared with large amounts of simulated data, simulated according to different dark matter models with the entire predicted population of low-mass haloes injected into the image. The likelihood of each dark matter model is then directly computed via comparison with the data. These methods fall in the category often described as approximate Bayesian computation or simulation based inference.
An obvious advantage is that the dark matter model is directly inferred, without the need to explicitly and expensively calculate the sensitivity. But this inference can be relied upon only as far as the assumptions made in producing the simulated data. This is also true for methods which forward model individual perturbers without gravitational imaging. In some cases these have been in good agreement with gravitational imaging (Nightingale et al., 2024). Other cases have produced false positive results, demonstrating the inherent risk of such a method (Hezaveh et al., 2016; Stacey et al., 2025).
However, forward-modelling methods have another significant advantage in that they can account for the effect of the entire population of perturbers, not just those large enough to be individually detectable with gravitational imaging. After all, in CDM, we expect hundreds or thousands of subhaloes larger than $10^{6}M_{\odot}$ to populate the haloes of lens galaxies. The sensitivity mapping necessary for dark matter inference using gravitational imaging does not currently account for this effect, if it exists. A correctly implemented forward-modelling method could detect or rule out such an effect, resulting in a more powerful inference on the dark matter model than considering individually detected objects only. Many of the works cited above include this effect implicitly, but do not show systematically that it exists, or measure its properties.
@@ 2 Method
We use the machine learning-based sensitivity mapping and substructure detection method introduced in O’Riordan et al. (2023, hereafter O’R23) and extended in O’Riordan & Vegetti (2024, hereafter O’R24). Our results are based on applying these models to different realisations of the same $100$ lens systems used in O’R24, which were selected to resemble SLACS and BELLS Einstein ring systems previously studied for substructure. In this work, we only consider the effect of subhaloes, i.e., member haloes of the main lens galaxy halo, and exclude haloes along the line of sight.
@@ 2.2 Model training
To train the machine learning models we simulate strong lens observations at HST-like resolution and depth. The lens galaxies are power-law ellipsoids and the source surface brightness distributions are taken from a de-noised sample of Hubble Deep Field galaxy observations. The light profile of the main lens galaxy, simulated as a single Sérsic profile, is subtracted up to its remaining Poisson noise. Different sources are used in training and testing to avoid overfitting.
@@
As in O’R24, we account for non-ellipticity in the lens mass distribution using multipole components. These are Fourier-like perturbations to the main lens density profile. To examine their effect on substructure detection, three versions of the model are trained. $\mathcal{M}_{1}$ includes no multipole perturbations in the training data. $\mathcal{M}_{2}$ includes them only up to amplitudes of $1$ per cent, and $\mathcal{M}_{3}$ includes them up to 3 per cent. $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$ use multipoles of orders $1$, 33, and $4$, with the amplitude of each order independently drawn from the model-dependent prior. $\mathcal{M}_{1}$ essentially represents most strong lensing models used until recently, a power-law ellipsoid plus external shear. $\mathcal{M}_{2}$ represents the best extension to $\mathcal{M}_{1}$, with non-elliptical components informed by the isophotes of elliptical galaxies, a model which is being used more commonly at present. $\mathcal{M}_{3}$ represents an extreme extension of $\mathcal{M}_{2}$, so that effects due to multipoles can be more easily borne out.
@@ 2.3 Sensitivity mapping
To create a sensitivity map, the models are shown realisations of the same strong lens system with a test mass substructure that moves across all pixels and over a range of masses. Each realised image is passed through the model and the probability of the positive class, i.e., that some amount of substructure is present, is recorded. A probability threshold is chosen, in this case $5\sigma$. The sensitivity map itself then follows in a straightforward way, as simply the lowest mass of subhalo in each pixel where the detection probability reached the threshold. We name this mass $m_{\mathrm{map}}$. We compute a sensitivity map for each lens in this way using a test mass range of $10^{7}<M_{\mathrm{max}}<10^{11}$, with 21 mass steps spaced log-uniformly. At each mass step, the $r_{\mathrm{max}}$ value of the subhalo is set to the median value from the $v_{\mathrm{max}}$-$r_{\mathrm{max}}$ relation of Moliné et al. (2022).
@@
Figure 3: Method to test the detectability of populations of individually undetectable substructures. A subhalo mass function, effectively the dark matter model, is chosen first, parametrised by $M_{\mathrm{hm}}$ and $f_{\mathrm{sub}}$. From this populations of subhaloes are drawn to produce images. The images are run through a subhalo detector CNN, which is the same used to create the sensitivity map for each lens. The sensitivity map and the subhalo mass function give $\mu_{\mathrm{sub}}$, the mean number of detectable objects, and in turn $p_{\mathrm{det}}$, the probability that the CNN makes a detection. The actual number of detections made by the CNN is $f_{\mathrm{det}}$. Comparing these quantities is the basis for the results in this paper.
@@ 2.4 Detecting a population of substructures
The number of detectable objects is Poisson distributed with mean $\mu_{\mathrm{sub}}$, so the probability that a given lens will have any detection, accounting for multiple individually detectable subhaloes, is
@@ 3.2 Excess as a function of subhalo mass
In region $(2)$, at $m_{\mathrm{cutoff}}\sim 10^{9.5}M_{\odot}$, the two curves begin to diverge. Figure˜2 shows that most pixels have no sensitivity below this mass. Accordingly, the predicted frequency of detections from the sensitivity map and mass function becomes flat as all newly added subhaloes are too small to be individually detected. However, the actual frequency of detections by the CNN increases until $m_{\mathrm{cutoff}}\sim 10^{7}M_{\odot}$. In regime $(3)$, below $m_{\mathrm{cutoff}}\sim 10^{7}M_{\odot}$, the excess of detections stops increasing. This shows that subhaloes in the intermediate mass range, i.e., in region $(2)$, have a detectable effect on the lensed images. Subhaloes at the lowest mass, i.e., for $m_{\mathrm{cutoff}}\lesssim 10^{7}M_{\odot}$ and below, do not have a detectable effect.
Figure 5: Properties of the subhaloes in the detectable population for different values of $M_{\mathrm{hm}}$. Upper frame: the modified mass function in eq.˜7 using the best fit values for our HST strong lenses. Lower frame: the ratio of the individual masses of the detectable population objects to the pixel sensitivities in our data, randomly drawn.
@@ 3.3 Accounting for the detectable population
The number $\mu_{\mathrm{pop}}$ can be thought of as the number of extra detectable objects if all detectable ‘populations’ were converted to equivalent individual objects. We calculate $\mu_{\mathrm{pop}}$ using a modified subhalo mass function which artificially increases the number of objects of mass $m$ by a factor $g(m)$, that is,
@@
This introduces three new parameters. The normalisation $n_{\mathrm{pop}}$ represents the strength of the effect in causing detections, and may be roughly interpreted as the ratio of the number of subhaloes in a detectable population to an equivalent individually detectable object. The mean and variance of $g(m)$ are $m_{\mathrm{pop}}$ and $\sigma_{\mathrm{pop}}^{2}$, but these are only weak indicators of the mass range of the detectable population. This is because of the shape of the mass function, with which $g(m)$ will be multiplied, where every decade lower in mass brings a decade more in number (see again fig.˜1). For example, if we find $m_{\mathrm{pop}}=10^{X}M_{\odot}$ with $\sigma_{\mathrm{pop}}=1$ (one order of magnitude in mass), much more than half of the population members will have mass $M_{\mathrm{max}}<10^{X}$.
@@
In the lower frame of fig.˜5 we illustrate the size of the population members compared to typical detectable individual subhaloes. We draw subhaloes from the pseudo-mass function used in eq.˜7 only and compare their masses with the sensitivity in a randomly drawn pixel from the $100$ mock lenses. For all values of $M_{\mathrm{hm}}$, the distribution peaks at $\sim 2$ orders of magnitude below the individually detectable mass. The absolute masses of this population (upper frame) depend on the data, so the peak in the detectable population mass function between $10^{8}M_{\odot}$ and $10^{9}M_{\odot}$ is specific to the BELLS and SLACS-like HST observations used here. However, it is very likely that the relative masses of the detectable population (lower frame) do not change with angular resolution. This is because sensitivity scales straightforwardly with angular resolution (Despali et al., 2022). Extrapolating the results here to higher angular resolutions would suggest that strong lens observations with e.g. VLBI, where objects as small as $10^{6}M_{\odot}$ are detectable, could be sensitive to populations of subhaloes whose constituent masses are $10^{5}M_{\odot}$ or lower. Replicating the results in this paper at that angular resolution is currently computationally intractable.
The use of $\mu_{\mathrm{pop}}$ here and its associated parameters is to correct the prediction for the rate of detections by the sensitivity map and the subhalo mass function, and to show that the observed excess in the rate of detections can be directly attributed to a low-mass halo population. It is useful for estimating the properties of the detectable population in this specific case where our machine learning models return a posterior probability only on the binary presence or absence of substructure. In a fully-forward modelling method (e.g. Brehmer et al., 2019) such a correction is not necessary, as the posterior on the dark matter model is already marginalised over all subhalo properties.
@@ 3.4 Improved constraints on dark matter
Here we consider the strength of constraint on the dark matter model, parametrised only by $M_{\mathrm{hm}}$, when the effect of the population either is or is not accounted for. We consider two simple instructive examples. In the first case, zero detections are made in our set of 100 lenses, which we label $\mathcal{D}_{0}$, and in the second case, 10 detections in 100 lenses, which we label $\mathcal{D}_{10}$. The case of 10 detections in 100 lenses represents an order of magnitude estimate of a realistic scenario given current substructure search efforts and the quality of the available data.
@@ 3.5 Angular structure in the lens galaxy
Using the same models from O’R24 we repeat the experiment shown in fig.˜4. This time, we use the two models described in section˜2.2. These models were trained with multipole perturbations in the lensing galaxy. This makes their substructure detections, and sensitivity maps, much more conservative.
@@ 4 Discussion
Here we discuss some limitations and implications of our work. We do not discuss aspects related to machine learning specific issues, e.g. quality of the training data, model architecture, etc. These issues were discussed for the same models used here in O’R23 and O’R24. In O’R24 a detailed comparison was made between the sensitivity maps produced by our machine learning models, and classical sensitivity maps produced by Despali et al. (2022). We found the overall structure of the sensitivity maps to be the same. That is, for a given lens and source configuration, the topology of the sensitivity maps are similar. However, we find in general that the machine learning method produces more conservative sensitivity maps. This is for two reasons. First, its much larger ‘prior’, i.e., its training data, includes all possible source and lens configurations, rather than the combinations which closely resemble the lens being modelled, as in the classical method. Second, and more importantly, is the extreme difficulty of the problem as presented to it, where almost all of the subhaloes presented in the training data and labelled as positive cases, are not detectable because they are too small or too far from the lensed images, or not concentrated enough. The result is an extremely cautious model where a $10\sigma$ detection in the classical method may be more like a $5\sigma$ detection in our case. In the context of this work, this conservative approach is actually an advantage, as it means the machine learning models never make spurious false positive detections unless a genuinely detectable subhalo signal is present in the data.
4.1 Systematic uncertainty due to lens galaxy properties